OpenStudy (anonymous):

determine weather each pair of equations represents perpendicular lines. show work y= -x+7 y -x =3

OpenStudy (anonymous):

those arent perpendicular lines, those are parallel because they have the same slope.

myininaya (myininaya):

right them both in y=mx+b form

myininaya (myininaya):

write*

OpenStudy (anonymous):

oops sorry i looked at the equations wrong ><

myininaya (myininaya):

y=-x+7 y=x+3 sp they are perpendicular since the slope of the first line is 1 and slope of the second line is -1 and their product is -1

OpenStudy (anonymous):

perpendicular lines have to have reciprocal slope. for example, y = 2x+2 y = -1/2x+5 those two are perpendicular lines because theyre negative reciprocals of eachother. see what i mean?

OpenStudy (anonymous):

yes i do see..thank u both

OpenStudy (anonymous):

np :)